# Attacking the Spike: On the Transferability and Security of Spiking Neural Networks to Adversarial Examples

### v3

@@ Abstract
Spiking neural networks (SNNs) have attracted much attention for their high energy efficiency and for recent advances in their classification performance. However, unlike traditional deep learning approaches, the analysis and study of the robustness of SNNs to adversarial examples remain relatively underdeveloped. In this work, we focus on advancing the adversarial attack side of SNNs and make three major contributions. First, we show that successful white-box adversarial attacks on SNNs are highly dependent on the underlying surrogate gradient technique, even in the case of adversarially trained SNNs. Second, using the best surrogate gradient technique, we analyze the transferability of adversarial attacks on SNNs and other state-of-the-art architectures like Vision Transformers (ViTs) and Big Transfer Convolutional Neural Networks (CNNs). We demonstrate that the adversarial examples created by non-SNN architectures are not misclassified often by SNNs. Third, due to the lack of an ubiquitous white-box attack that is effective across both the SNN and CNN/ViT domains, we develop a new white-box attack, the Auto Self-Attention Gradient Attack (Auto-SAGA). Our novel attack generates adversarial examples capable of fooling both SNN and non-SNN models simultaneously. Auto-SAGA is as much as $91.1\%$ more effective on SNN/ViT model ensembles and provides a 3×3$\times$3 × boost in attack effectiveness on adversarially trained SNN ensembles compared to conventional white-box attacks like Auto-PGD. Our experiments and analyses are broad and rigorous covering three datasets (CIFAR-10, CIFAR-100 and ImageNet), five different white-box attacks and nineteen classifier models (seven for each CIFAR dataset and five models for ImageNet).
Introduction
@@
On the other hand, the vulnerability of deep learning models to adversarial examples  (Goodfellow et al. 2014) is one of the main topics that has received much attention in recent research. An adversarial example is an input that has been manipulated with a small amount of noise such that a human being can correctly classify it. However, the adversarial example is misclassified by a machine learning model with high confidence. A large body of literature has been devoted to the development of both adversarial attacks (Tramer et al. 2020) and defenses (Zhang et al. 2020a) for CNNs.
As SNNs become more accurate and more widely adopted, their security vulnerabilities will emerge as an important issue. Recent work has been done to study some of the security aspects of the SNN (El-Allami et al. 2021; Sharmin et al. 2019, 2020; Goodfellow et al. 2015; Kundu et al. 2021; Liang et al. 2021), although not to the same extent as CNNs. To the best of our knowledge, there has not been rigorous analyses done on how the different choice of gradient estimations can effect white-box SNN attacks. In addition, it is an open question whether SNN adversarial examples are misclassified by other state-of-the-art models like Vision Transformers. Finally, there has not been any general attack method developed to break both SNNs and CNNs/ViTs simultaneously. Thus in our paper, we specifically focus on three key adversarial aspects:
@@
    How are white-box attacks on SNNs affected by different SNN surrogate gradient estimation techniques?
@@
    As SNNs have shown more robustness in previous studies (Sharmin et al. 2019; Liang et al. 2021), do adversarial examples generated by SNNs transfer to other models such as Vision Transformers and CNNs and vice versa?
@@
    Are there a white-box attacks that can effectively target both SNNs and CNNs/Vision Transformers, closing the transferability gap and achieving a high success rate?
These three questions are intrinsically linked: by first focusing on the surrogate gradient, we can develop effective SNN white-box attacks. These precise SNN attacks are the foundation for analyzing the transferability of adversarial examples between SNNs and other SOTA architectures (as posed in our second question). Based on the outcome of the second question (low attack transferability, which we show in Section SNN Transferability Study), we further develop a new attack capable of breaking SNN and non-SNN models simultaneously.
We organize the rest of our paper as follows: in Section SNN Models, we briefly introduce the different types of SNNs, whose security we analyze. In Section SNN Surrogate Gradient Estimation, we analyze the effect of seven different surrogate gradient estimators on white-box adversarial attacks. We show that the choice of surrogate gradient estimator is highly influential and must be carefully selected. In Section SNN Transferability Study, we use the best surrogate gradient estimator to study the transferability of adversarial examples generated by SNNs. Our transferability experiments demonstrate that traditional white-box attacks are not effective on both SNNs and other non-SNN models simultaneously, mandating the need for a new multi-model attack. In Section Transferability and Multi-Model Attacks, we develop a new multi-model attack capable of creating adversarial examples that are misclassified by both SNN and non-SNN models. Our attack is called the Auto Self-Attention Gradient Attack (Auto-SAGA) and we empirically show its superiority to MIM (Dong et al. 2018), PGD (Madry et al. 2018) and Auto-PGD (Croce et al. 2020).
@@
Overall, we conduct rigorous analyses and experiments with 19 models across three datasets (CIFAR-10, CIFAR-100 and ImageNet) and four adversarial training methods. Our surrogate gradient and transferability results yield new insights into SNN security. Our newly proposed attack (Auto-SAGA) works on SNN/ViT/CNN ensembles with higher attack success rates (as much as $91.9\%$ improvement) and boosts the attack effectiveness on adversarially trained SNN ensembles by 3×3$\times$3 × over conventional white-box attacks like Auto-PGD. All three of our major paper contributions significantly advance the security development of SNN adversarial machine learning.
SNN Models
@@
In this section, we discuss the basics of the SNN architecture and of neural encoding. Widely used Leaky Integrate and Fire (LIF) neuron can be described by a system of difference equations as follows (Shrestha et al. 2022):
@@
where $V[t]$ denotes neuron’s membrane potential. $\alpha\in(0,1]$ is a time constant, which controls the decay speed of membrane potential. When $\alpha=1$, the model becomes Integrate and Fire (IF) neuron. $S_{i}[t]$ and $w_{i}$ are $i_{th}$ input and the associated weight. $\vartheta$ is the neuron’s threshold, $O[t]$ is the neuron’s output function, $u(\cdot)$ is the Heaviside step function. If $V[t]$ exceeds the threshold $\vartheta$, the neuron will fire a spike, hence $O[t]$ will be 1. Then, at the next time step, $V[t]$ will be decreased by $\vartheta$ in a procedure referred to as a reset (Shrestha et al. 2022).
@@
Note that, in contrast to the continuous input domains of DNNs, in SNNs information is represented by discrete, binary spike trains. Therefore, data has to be mapped to spike domain for an SNN to process, this procedure is referred to as neural encoding, which plays an essential role in high performance SNN applications (Shrestha et al. 2022). A popular way to achieve such mapping is the direct encoding (Wu et al. 2019; Rathi et al. 2021). This method uses Current-Based neuron (CUBA) as the first layer in an SNN’s architecture. CUBA accepts continuous values instead of spikes (Fang et al. 2020a) such that a pixel value can be converted into a spike train and then directly fed to the SNN. Because direct encoding can reduce inference latency by a factor of $5{-}100$ (Rathi et al. 2021), recent works have achieved state-of-the-art results with this coding scheme (Rathi et al. 2021; Fang et al. 2021a; Kundu et al. 2021; Fang et al. 2020b), all experiments employ direct coding in this work.
SNN Training
Spiking neuron’s non-differentiable activation makes directly applying backpropagation (BP) difficult (Zhang et al. 2020b; Tavanaei et al. 2019). Training an SNN requires different approaches, which can be categorized as follows:
@@
Conversion-based training. A common practice is to use spike numbers in a fixed time window (spike rate) to represent a numerical value. The input strength and neuron’s spike rate roughly have a linear relation (such behavior is similar to ReLU function in DNNs (Diehl et al. 2015)). Therefore, it is possible to pretrain a DNN model and map the weights to an SNN. However, simply mapping the weights suffers from performance degradation due to non-ideal input-spike rate linearity, over activation, and under activation (Rathi et al. 2021; Diehl et al. 2015). Additional post-processing and fine tuning are required to compensate for the performance degradationu such as weight-threshold balancing (Diehl et al. 2015).

### v4

# Attacking the Spike: On the Security of Spiking Neural Networks to Adversarial Examples

@@ Abstract
Spiking neural networks (SNNs) have attracted much attention for their high energy efficiency and for recent advances in their classification performance. However, unlike traditional deep learning approaches, the analysis and study of the robustness of SNNs to adversarial examples remain relatively underdeveloped. In this work, we focus on advancing the adversarial attack side of SNNs and make three major contributions. First, we show that successful white-box adversarial attacks on SNNs are highly dependent on the underlying surrogate gradient estimation technique, even in the case of adversarially trained SNNs. Second, using the best single surrogate gradient estimation technique, we analyze the transferability of adversarial attacks on SNNs and other state-of-the-art architectures like Vision Transformers (ViTs), as well as CNNs. Our analyzes reveal two key areas where SNN adversarial attacks can be enhanced: no white-box attack effectively exploits the use of multiple surrogate gradient estimators for SNNs, and no single model attack is effective at generating adversarial examples misclassified by both SNNs and non-SNN models simultaneously.
For our third contribution, we develop a new attack, the Mixed Dynamic Spiking Estimation (MDSE) attack to address these issues. MDSE utilizes a dynamic gradient estimation scheme to fully exploit multiple surrogate gradient estimator functions. In addition, our novel attack generates adversarial examples capable of fooling both SNN and non-SNN models simultaneously. The MDSE attack is as much as $91.4\%$ more effective on SNN/ViT model ensembles and provides a $3\times$ boost in attack effectiveness on adversarially trained SNN ensembles, compared to conventional white-box attacks like Auto-PGD. Our experiments are broad and rigorous, covering three datasets (CIFAR-10, CIFAR-100 and ImageNet) and nineteen classifier models (seven for each CIFAR dataset and five models for ImageNet). We will release a full publicly available code repository for the models and attacks upon publication.
keywords:
Spiking Neural Network , SNN , Machine Learning Security , Adversary Attack , Attack Transferability
equalequalfootnotetext: These authors contributed equally to this work.
@@
[…]
1 Introduction
[…]
On the other hand, the vulnerability of deep learning models to adversarial examples  [12] is one of the main topics that has received much attention in recent research. An adversarial example is an input that has been manipulated with a small amount of noise such that a human being can correctly classify it. However, the adversarial example is misclassified by a machine learning model with high confidence. A large body of literature has been devoted to the development of both adversarial attacks [13] and defenses [14] for CNNs.
As SNNs become more accurate and more widely adopted, their security vulnerabilities will emerge as an important issue. Recent work has been done to study some of the security aspects of the SNN [15, 16, 17, 18, 19, 20], although not to the same extent as CNNs. A unique challenge arises in the study of SNN due to spiking neuron’s non-differentiable binary activation, i.e., neuron’s output can only be 1 (fire) or 0 (not fire). As true gradients do not exist in SNNs, surrogate gradient [21], which is a technique to estimate the gradient of spikes, has been proposed to enable BP. White-box attacks in adversarial machine learning rely on accurate gradient calculation. However, the choice of surrogate estimator for the gradient calculations in SNNs is highly flexible. How the gradient estimation affects white-box remains unknown. To the best of our knowledge, there have not been rigorous analyses done on how the different choices of gradient estimations can effect white-box SNN attacks. In addition, it is an open question whether SNN adversarial examples are misclassified by other state-of-the-art models like Vision Transformers (ViTs). Finally, there has not been any general attack method developed to break both SNNs and CNNs/ViTs simultaneously. Thus in our paper, we specifically focus on three key security aspects:
@@ 1 Introduction
    How do the use of different SNN gradient estimation functions impact the effectiveness of white-box attacks?
@@
    As SNNs have shown more robustness in previous studies [16, 20], do adversarial examples generated by SNNs transfer to other models such as Vision Transformers and CNNs and vice versa?
@@
    Can white-box attacks leverage different gradient estimation functions to more effectively attack SNNs? Additionally, can a white-box attack be developed that effectively target both SNNs and CNNs/Vision Transformers, closing the transferability gap and achieving a high success rate?
Paper Organization: These three questions are intrinsically linked and form the outline for our paper. After introducing the types of SNNs in Section 2.4 we show the choice of gradient estimator plays a major role in the success of SNN white-box attacks in Section 3. Then, using the single best gradient estimator, we analyze the transferability of adversarial examples between SNNs and other SOTA architectures (as posed in our second question) in Section 4. Based on the outcome of the second question (low attack transferability), an important attack issue arises: current SOTA white-box attacks cannot break an ensemble of SNNs and non-SNN models. To solve this issue, we further develop a new attack, the Mixed Dynamic Spiking Estimation (MDSE) attack in Section 5. The advantages of our new attack are two-fold: Through dynamic spiking surrogate gradient estimation we create a more effective SNN specific attack framework. Second, by mixing gradients from multiple models we are able to craft adversarial examples that are misclassified by both non-SNN and SNN models simultaneously, bridging the transferability gap. We empirically demonstrate the superiority of the MDSE attack to MIM [22], PGD [23], SAGA [24] and Auto-PGD [25] in Section 6.
@@
Main Contributions: Overall, we conduct rigorous analyses and experiments with 19 models across three datasets (CIFAR-10, CIFAR-100, and ImageNet) and four adversarial training methods. We consider two recently proposed SNN-based adversarial training methods: Temporal Information Concentration (TIC) [26] and HIRE [19]. Additionally, we explore SNN models trained with techniques originally designed for CNNs, such as Diffusion Model (DM) enhanced adversarial training [27] and Friendly Adversarial Training (FAT) [14]. Our surrogate gradient estimator results on normal and adversarially trained SNNs consistently show that an optimal SG is crucial for accurately evaluating the robustness of SNNs. The transferability results highlight the low attack transferability among SNNs and non-SNNs, providing new insights into SNN security. Our newly proposed attack, MDSE, achieves higher attack success rates on SNN/ViT/CNN ensembles, with improvements of up to $91.4\%$. Additionally, MDSE is three times more effective than conventional white-box attacks like Auto-PGD when targeting adversarially trained SNN ensembles. These findings significantly advance the security development of SNN adversarial machine learning.
2 Related Work and Spiking Neural Network Background
2.1 Related Work in Spiking Neural Network Security
The advent of the Fast Gradient Sign Method (FGSM) [18] in 2015 and other attacks [28] in 2013 spurred new security questions about machine learning models. Collectively, the security problems relating to machine learning models has been termed adversarial machine learning [28]. In this work, we specifically focus on the white-box evasion attack where an adversary manipulates the input to a machine learning model to cause misclassification.
Different security aspects of Spiking Neural Networks (SNNs) have been studied in the field of adversarial machine learning. In [29] the privacy issues related to data leakage during SNN training was studied. Backdoor attacks against SNNs were developed in [30] and the vulnerability of SNNs to membership inference attacks was explored in [31]. Evasion attacks against SNNs were first analyzed in [32, 15, 17]. While significant work has also been published to robustify SNNs to evasion attacks, in general, it has been shown that adversarial training is one of the only defense techniques to withstand the test of time. This is due to the fact that many other defenses have been broken by adaptive attacks [13, 33, 34]. Hence in this work for the SNN defenses, we focus on the different adversarial training regimes that have been developed [26, 19, 14, 27]. The adversarial defense techniques and our reasons for selecting them are further detailed in subsection 2.6. In the next subsection we briefly introduce adversarial machine learning evasion attacks.
2.2 Related Work: Challenges of Spike-based Input
Adversarial attacks to SNNs presents varying challenges depending on the input data type. SNNs can process both regular datasets (e.g., MNIST, CIFAR-10) and spike-based neuromorphic datasets (e.g. N-MNIST, DVS gesture). For regular datasets, which are defined in in continuous domain $\mathbb{R}$, conventional image inputs can be converted into spike trains [35, 7]. And with surrogate gradient, adversarial perturbations can be computed in the continuous domain and directly applied to input images. However, neuromorphic datasets represent information using binary spike trains $\in\{0,1\}$, where each value indicates either the presence or absence of a spike. This fundamental difference in data representation makes conventional adversarial attack methods incompatible with spike-based inputs, as adding continuous-valued perturbation to spikes would violate the binary nature. There are works proposed specialized techniques to address this challenge  [20, 36]. Combining these approaches with our method is possible but beyond our current scope. Therefore, in this work, we focus on regular datasets, leaving the integration with spike-based attack methods for future investigation.
2.3 Related Work: Evasion Attacks
Fast Gradient Sign Method (FGSM): the Fast Gradient Sign Method (FGSM) [18] is a white-box attack that generates adversarial examples by adding noise to the clean image in the direction of the gradients of the loss function:
$x_{\text{adv}}=x+\epsilon\cdot\text{sign}(\nabla_{x}\mathcal{L}(x,y;\theta))$ (1)
where $x_{\text{adv}}$ is the adversarial example, $x$ is the original input, $\epsilon$ is the attack step size parameter, $\mathcal{L}$ is the loss function of the targeted model, $y$ is the true label, $\theta$ represents the model parameters, and $\nabla_{x}$ denotes the gradient with respect to the input $x$. The attack performs only a single step of perturbation size $\epsilon$, and applies noise in the direction of the sign of the gradient of the model’s loss function.
Projected Gradient Descent: the Projected Gradient Descent attack (PGD) [23] is a modified version of the FGSM attack that implements multiple attack steps. The attack attempts to find the minimum perturbation, bounded by $\epsilon$, which maximizes the model’s loss for a particular input, $x$. In the randomized version of the attack a random perturbation is generated on a ball centered at $x$ and with radius $\epsilon$. Adding this noise to $x$ gives the initial adversarial input, $x_{0}$. From here, the attack begins an iterative process that runs for $k$ steps. During the $i^{th}$ attack step the perturbed image, $x_{i}$, is updated as follows:
$x_{i}=P(x_{i-1}+\alpha\cdot\text{sign}(\nabla_{x_{i-1}}\mathcal{L}(x_{i-1},y;\theta))$ (2)
Momentum Iterative Method: the Momentum Iterative Method (MIM) [22] applies momentum techniques seen in machine learning training to the domain of adversarial machine learning. Similar to those learning methods, the MIM attack’s momentum allows it to overcome local minima and maxima. The attack’s main formulation is similar to the formulation seen in the PGD attack. Each attack iteration is calculated as follows:
$x_{i}=\text{clip}(x_{i-1}+\frac{\epsilon}{t}\cdot\text{sign}(g_{i}))$ (3)
where $x_{i}$ represents the adversarial input at iteration $i$, clip is a clipping operation to keep the adversarial example within the valid image values, $\epsilon$ is the total attack magnitude, and $t$ is the total number of attack iterations. $g_{i}$ represents the accumulated gradient at step $i$ and is calculated as follows:
[…]
where $\mu$ represents a momentum decay factor. The MIM attack has been shown to create adversarial examples that are transferable [24], in the sense that multiple models misclassify MIM adversarial examples.
Auto-Projected Gradient Descent: the Auto-PGD [25] attack is a budget-aware, step size-free variant of PGD. The algorithm partitions the available $N_{\text{iter}}$ iterations to first search for a good initial point. Then, in the exploitation phase, it progressively reduces the step size to maximize the attack results. However, the reduction in step size is not pre-determined but is governed by the optimization trend: if the loss value grows sufficiently fast, then the step size is likely appropriate and no change is made. Otherwise, Auto-PGD reduces the step size in an attempt to improve the attack algorithm performance. The gradient update of Auto-PGD follows closely the classic algorithm and adds a momentum term. Let $\eta_{i}$ be the step size at iteration $i$, then the update step is as follows:
$z_{i+1}$ $=P\left(x_{i}+\eta_{i}\nabla f(x_{i})\right)$ (5)
$x_{i+1}$ $=P\left(x_{i}+\alpha\cdot(z_{i+1}-x_{i})+(1-\alpha)\cdot(x_{i}-x_{i-1})\right)$
where $\alpha\in[0,1]$ regulates the influence of the previous update on the current one, $z_{i}$ is the intermediate perturbed point, and $P$ is the projection operation.
2.4 Spiking Neural Network Types
@@
In this subsection, we discuss the basics of the SNN architecture and of neural encoding. Widely used Leaky Integrate and Fire (LIF) neuron can be described by a system of difference equations as follows [2]:
@@
where $V[t]$ denotes neuron’s membrane potential. $\alpha\in(0,1]$ is a time constant, which controls the decay speed of membrane potential. When $\alpha=1$, the model becomes Integrate and Fire (IF) neuron. $S_{i}[t]$ and $w_{i}$ are $i_{th}$ input and the associated weight. $\vartheta$ is the neuron’s threshold, $O[t]$ is the neuron’s output function, $u(\cdot)$ is the Heaviside step function. If $V[t]$ exceeds the threshold $\vartheta$, the neuron will fire a spike, hence $O[t]$ will be 1. Then, at the next time step, $V[t]$ will be decreased by $\vartheta$ in a procedure referred to as a reset [2].
@@
Note that, in contrast to the continuous input domains of DNNs, in SNNs information is represented by discrete, binary spike trains. Therefore, data has to be mapped to the spike domain for an SNN to process, which is known as neural encoding [2]. A popular way to achieve such a mapping is by using direct encoding [35, 7]. This encoding can reduce inference latency by a factor of 51005{-}100 [7]. Recent works have achieved state-of-the-art results with this coding scheme [7, 10, 19, 9]. Hence, all experiments in this paper employ direct coding.
2.5 Spiking Neural Network Training
@@
The neurons within the SNN have non-differentiable activation functions, which makes directly applying BP challenging [37, 38]. Broadly speaking, there are two common techniques for training an SNN, conversion based or surrogate gradient based training. In conversion based training, it is possible to pre-train a DNN model and map the weights to an SNN. However, simply mapping the weights suffers from performance degradation due to non-ideal input-spike rate linearity, over activation, and under activation [7, 39]. Additional post-processing and fine tuning are required to compensate for the performance degradation such as weight-threshold balancing [39].
A second way to train SNNs is through the use of surrogate gradient BP. Equation 6a - 6c reveal that SNNs have a similar form to Recurrent Neural Networks (RNNs). The membrane potential is dependent on input and historical states. Equation 6a is actually differentiable, thereby making it possible to unfold the SNN and use BP to train it.
Following the standard way to unfold the SNN, we can derive the Backpropagation rule. Let $o_{i}^{l}[t]$ represents the output of $i_{th}$ neuron in layer $l$ at time $t$, and it derivative with respect to the loss function $\mathcal{L}$ be $\delta^{l}_{i}[t]=\frac{\partial\mathcal{L}}{o^{l}_{i}[t]}$. By applying the chain rule, $\delta^{l}_{i}[t]$ can be calculated recursively[21]:
$\delta^{l}_{i}[t]=u^{\prime}(V^{l}_{i}[t])(\sum\delta^{l+1}_{n})W^{\intercal,l}_{i}k+\delta^{l}_{i}[t+1]\vartheta)$ (7)
where $u^{\prime}(\cdot)$ is the derivative of function $u(\cdot)$, which will be discussed in section 3. Weight can be updated as [21]:
$\Delta W^{l}_{ij}\propto\frac{\partial\mathcal{L}}{W^{l}_{ij}}=\sum_{t=0}^{T-1}\delta^{l}_{i}[t]o^{l-1}_{j}[t]$ (8)
The challenge is Equation 6c, i.e. the Heaviside step function $u(\cdot)$ is non-differentiable. To overcome this issue, the surrogate gradient method has been proposed [21], which allows the Heaviside step function’s derivative $u^{\prime}(\cdot)$ to be approximated by some smooth function. Using a surrogate gradient enables SNN training with BP and achieves comparable performance to DNNs [8, 10]. There are multiple viable choices for the surrogate gradient method.
2.6 Spiking Neural Network Adversarial Training
We also consider defenses based on SNNs, in addition to vanilla (undefended) SNN models. One of the most common ways to defend against adversarial attacks is through adversarial training [23]. In our work, we consider two recently proposed SNN-based adversarial training methods: Temporal Information Concentration (TIC) [26] and HIRE [19]. Additionally, we explore SNN models trained with techniques originally designed for CNNs, such as Diffusion Model (DM) enhanced adversarial training [27] and Friendly Adversarial Training (FAT) [14].
Temporal Information Concentration (TIC) [26] indicates the information in SNN shifts from latter timesteps to earlier timesteps as training progresses. The defense proposed a loss function to control the Fisher information value at each timestep:
$L_{t}(\theta,\alpha)=|L_{t}(\theta)-\alpha|$ (9)
$L(\theta,\alpha)=\frac{1}{T}\sum_{t=1}^{T}L_{t}(\theta,\alpha)$ (10)
We apply Eq. 9 across $T$ timesteps to force the loss function to a value around $\alpha$, ensuring that the Fisher information shows a similar trend for all timesteps. The key takeaway is that the SNN model exhibiting temporal concentration behavior (smaller Fisher trace as time goes on) might have better robustness.
Why we selected it: TIC was selected for its potential to specifically address the unique temporal dynamics of SNNs with improved robustness that is directly related to the SNN features.
HIRE-SNN (Spike Timing Dependent Backpropagation) [19] is a training algorithm designed to enhance the inherent robustness of conversion-based SNNs. It partitions the total time steps $T$ into $N$ equal-length periods. The gradients with respect to the weights $\delta_{w}$ and perturbations $\kappa$, as well as threshold $v_{t}$ and leak $l_{k}$ parameters, are calculated and updated over small intervals of $\left\lfloor\frac{T}{N}\right\rfloor$ steps. The HIRE-SNN training process involves calculating gradients and updating weights based on the following steps:
$\delta_{w}\leftarrow\mathbb{E}_{(x,y)\in{B}}\left[\nabla_{w}\mathcal{L}\left(g(x+\kappa,y;\frac{T}{N})\right)\right]$ (11)
$\delta_{x}\leftarrow\left[\nabla_{x}\mathcal{L}\left(g(x+\kappa,y;\frac{T}{N})\right)\right]$ (12)
$\kappa\leftarrow\text{clip}(\kappa+\epsilon_{s}\cdot\text{sign}(\delta_{x}),-\epsilon_{t},\epsilon_{t})$ (13)
$\mathbf{W}\leftarrow\mathbf{W}-\eta\cdot\delta_{w}$ (14)
In these equations, $(x,y)$ represents the input data and label, B{B} denotes the batch of data, and $\delta_{x}$ represents the gradients with respect to inputs. $\mathbf{W}$ represents the model weights. $\epsilon_{s}$ is the step size for the perturbation, $\epsilon_{t}$ is the perturbation limit, $\eta$ is the learning rate. This method allows the model to be trained with various adversarial image variants without incurring additional training time.
Why we selected it: HIRE was chosen for its ability to enhance SNN robustness against temporal perturbations, offering a strong benchmark for assessing the effectiveness of our proposed adversarial attacks across different SNN models.
Friendly Adversarial Training (FAT) [14] focuses on identifying the least adversarial data that minimizes the loss among the adversarial data that is misclassified. This training approach employs a modified version of PGD called PGD-K-$\tau$. In PGD-K-$\tau$, $K$ refers to the number of iterations used for PGD, and $\tau$ is a hyperparameter that allows early stopping in the PGD generation of adversarial examples if the sample is already misclassified. The FAT method involves updating the model parameters $\theta$ as follows:
$\theta\leftarrow\theta-\eta\frac{1}{m}\sum_{i=1}^{m}\nabla_{\theta}\ell(f_{\theta}(\tilde{x}_{i}),y_{i})$ (15)
where $\theta$ represents the model parameters, $\tilde{x}_{i}$ represents the adversarially perturbed input, $y_{i}$ is the true label, $\ell$ is the loss function, $\eta$ is the learning rate, and $m$ is the batch size.
Why we selected it: FAT is an adversarial training method that can maintain higher clean accuracy due to its early stopping PGD algorithm during training. However, FAT has only been tested with CNN variants and never with SNN models. As FAT is one prominent recent adversarial training algorithm, testing its effectiveness with SNNs is of interest.
Diffusion Model (DM) Enhanced Adversarial Training [27] involves the use of class-conditional elucidating diffusion models (EDM) [40] to generate augmented datasets for CIFAR-10 and CIFAR-100. These datasets are used in the TRadeoff-inspired Adversarial DEfense via Surrogate-loss minimization (TRADES) [41] pipeline, which employs a classification-calibrated loss theory to balance accuracy and robustness. The loss function used in TRADES is:
$\mathcal{L}_{\text{TRADES}}=\mathcal{L}_{\text{CE}}(f_{\theta}(x),y)+\beta\cdot\max_{x^{\prime}\in\mathcal{B}(x,\epsilon)}\mathcal{L}_{\text{KL}}(f_{\theta}(x),f_{\theta}(x^{\prime}))$ (16)
where $\mathcal{L}_{\text{CE}}$ is the cross-entropy loss, $f_{\theta}(x)$ is the model output for input $x$ with parameters $\theta$, $y$ is the true label, $\beta$ is the hyperparameter to control the trade-off, and $\mathcal{L}_{\text{KL}}$ is the Kullback-Leibler divergence between the outputs of the original input $x$ and the perturbed input $x^{\prime}$. $\mathcal{B}_{p}(x,\epsilon):=\{x^{\prime}\mid\|x^{\prime}-x\|_{p}\leq\epsilon\}$ denotes that the input $x^{\prime}$ is constrained into the $\ell_{p}$ norm, where $\epsilon$ is the maximum perturbation constraint.
Why we selected it: DM was included to assess how data augmentation and loss optimization techniques originally developed for CNNs perform in enhancing SNN robustness against adverarial attacks. In addition, DM provides SOTA robustness results on CIFAR-10 with CNN architectures, making it an ideal candidate to implement and test in the SNN domain.
The Importance of Adversarial Training: To the best of our knowledge, we are the first to implement DM and FAT adversarial training techniques on SNNs. We are also the first to compare existing SNN-specific adversarial training (TIC and HIRE) to DM and FAT. While this alone is not a major contribution, in the context of developing adversarial attacks, it is critical to include these types of analyses. This is because existing adversarial attacks can readily be adapted to new undefended architectures [24] yielding a high attack success rate. However, on defended models or model ensembles, existing adversarial attacks may not be effective. Experimenting with adversarial training defense methods are key for accurately assessing the robustness of SNNs to SOTA adversarial attacks. Many other adversarial training algorithms exist and are being proposed or developed recently [42, 43]. While we cannot test all of them, we do include [43] for discussion and comparison in Section 6.
3 Surrogate Gradient Estimation
In both neural network training and white-box adversarial machine learning attacks, the fundamental computation requires backpropagating through the model. Due to the non-differentiable structure of SNNs [21], this requires using a surrogate gradient estimator. In [44], it was shown that gradient based SNN training was robust to different derivative shapes. In [35], it was demonstrated that there are multiple different gradient estimators that can lead to reasonably good performance on MNIST, N-MNIST and CIFAR-10.
While there exist multiple viable surrogate gradient estimators for SNN training, in the field of adversarial machine learning, precise gradient calculations are paramount. Incorrect gradient estimation on models leads to a phenomenon known as gradient masking [45]. Models that suffer from gradient masking appear robust, but only because the model gradient is incorrectly calculated in white-box attacks performed against them. This issue has led to many published models and defenses to claim security, only to later be broken when correct gradient estimators were implemented [13]. To the best of our knowledge, this issue has not been thoroughly explored for SNNs in the context of adversarial examples. Hence, we run white-box attacks on SNNs using different surrogate gradient estimators, to empirically understand their effect on attack success rate. In our analyses, we experiment with undefended SNNs and four types of adversarial trained SNN models.
3.1 Representative Surrogate Gradient Functions
Figure 1: Different surrogate gradient functions.
Here we give a brief introduction to some representative surrogate gradient functions. We denote the derivative of Heaviside Step Function $u(x)$ as $u^{\prime}(x)$, threshold as $\vartheta$. The surrogate gradients investigated in this work are discussed as follows, and their shapes are shown in Figure 1.
Sigmoid [46] is a pioneer work which studies gradient estimation of non-smooth neuron. It indicates that a hard threshold function’s derivative can be approximated by that of a Sigmoid function. Such that $u^{\prime}(x)$ can be approximated as:
$u^{\prime}(x)\approx\frac{e^{\vartheta-x}}{(1+e^{\vartheta-x})^{2}}$ (17)
Erfc [9] takes a bio-inspired approach, it proposes to use the Poisson neuron’s spike rate function, which can be characterized by a complementary error function (erfc). Its derivative is given in (18), where $\sigma$ controls the sharpness.
Arctan [11, 10] used Arctangent function as surrogate gradient, achieving state-of-the-art results on various datasets. The surrogate gradient is given by:
$u^{\prime}(x)\approx\frac{1}{1+\pi^{2}(x-\vartheta)^{2}}$} (19)
Piece-wise linear function (PWL) [21] is the first work that formally established the framework of Surrogate Gradient method. It studied PWL function as gradient surrogate. In addition, PWL is also used in [7, 47]. Its formulation is given by:
$u^{\prime}(x)\approx\max(0,\vartheta-|x|)$ (20)
Fast Sigmoid [48] uses Fast Sigmoid as a replacement of the Sigmoid function, the purpose is to avoid expensive exponential operation and to speed up computation. It is defined as:
$u^{\prime}(x)\approx\frac{1}{1+(1+|x-\vartheta|)^{2}}$ (21)
Piece-wise Exponential [8] suggests that Probability Density Function (PDF) for a spiking neuron to change its state (fire or not) can approximate the derivative of the spike function. Spike Escape Rate, which is a piece-wise exponential function, can be a good candidate to characterize this probability density. It is given by (22), where $\alpha$ and $\beta$ are two hyperparamaters.
$u^{\prime}(x)\approx\frac{1}{\alpha e^{-\beta|x-\vartheta|}}$ (22)
Rectangular function is used by [49, 35], which are two representative works that empirically demonstrated that Surrogate Gradient together with Backpropagation Through Time can be used to train high performance SNNs. It is given by (23), where $\alpha$ is a hyperparameter that controls height and width.